Whats the frequency of green light waves with a wavelength of 520 Nm? What fromula would I use?
Answer #1

The formula relating the frequency f and wavelength λ is

fλ=c or f=c/λ

where c is the speed of light.

In your case, λ = 520nm = 520 × 10⁻⁹ m and c = 2.998 × 10⁸ m·s⁻¹.

∴ f=c/λ

=2.998×10⁸m⋅s⁻¹/520×10⁻⁹m=5.765 × 10¹⁴ s⁻¹